40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Feb 3 14:44:33 2012, Jenne, Update, IOO, MC SUS misalignment SUSjump_3Feb2012.png
    Reply  Fri Feb 3 15:49:10 2012, Jenne, Update, IOO, MC SUS misalignment 
    Reply  Fri Feb 3 17:17:47 2012, Den, Update, IOO, MC SUS misalignment 
Message ID: 6246     Entry time: Fri Feb 3 15:49:10 2012     In reply to: 6244
Author: Jenne 
Type: Update 
Category: IOO 
Subject: MC SUS misalignment 

[Jenne, Den]

We moved the MC approximately back to where the sensors for each optic used to be (mostly touching MC2, but a little bit of MC1 to help the refl get back to its max value).  MC is now locked, and with the help of the WFS it's back to nominal.  I forgot to disable the WFS, so I think we aren't perfectly aligned, but we're close enough for the WFS to get us the rest of the way.  We're heading over to JClub right now, so we're going to leave it as-is.

ELOG V3.1.3-